Find three solutions to y=-x+2. Use x=0​, x=2​, and x=-9.
Answer by mananth(16946)

y=-x+2. Use x=0​, x=2​, and x=-9.
y=-x+2
When x=0, y=2
when x=2
y=-2+2 =0
x=2,y=0
you try the third one